Matano upon retirement came across the following investment regimes;

a) 12% p.a compounded annually

b) 11% p.a compounded quarterly

c) 10% p.a. compounded monthly

d) 9.85% p.a compounded continuously

 

advise

i) As a rational lender which one would you choose?                                               (3 marks)

ii) As a rational borrower which of the regimes would you choose?                         (3 marks)

 


1
Expert's answer
2021-06-20T17:58:04-0400

"A=P(1+\frac{r}{n})^{nt}"

Interest=A-P

where A is the amount ,P is the initial principal balance, r is the interest rate, n is the number of times interest is compounded per time period and t is the number of time periods

For example let's take P=2000, t=4years

(a)

"A = 2,000.00(1 + \frac{0.12}{1})^{(1)(4)}\\
A = 2,000.00(1 + 0.12)^4\\
A = \$3,147.04\\
I=3147.04-2000=\$1147.04"


(b)

"A = 2,000.00(1 + 0.11/4)^{4×4}\\
A = 2,000.00(1 + 0.0275)^{16}\\
A =\$3,087.02\\
I=3087.02-2000=1087.02"


(C)

"A = 2,000.00(1 + 0.1/12)^{12×4}\\
A = 2,000.00(1 + 0.008333333)^{48}\\
A = \$2,978.71\\
I=2978.71-2000=\$978.71"


(d)

"A = 2,000.00(1 + 0.0985/365){365×4}\\
A = 2,000.00(1 + 0.000269863)^{1460}\\
A = \$2,965.64\\
I=2965.64-2000=965.64"


(i)A rational lender would choose option a because it earns highest interest.

(ii)A rational borrower should choose option d because it earns low interest rate.